OpenStudy (anonymous):

The simple interest on an investment of $8000 for one year is $360. What is the interest rate? A. 0.045% B. 0.45% C. 4.5% D. 45%

OpenStudy (anonymous):

probably c but I'm not 100% sure

OpenStudy (welshfella):

I = PRT/100 where P = principal (8000), R = rate and T = time (1) I = 360

OpenStudy (welshfella):

sove for R 360 = 8000R / 100

OpenStudy (welshfella):

* solve

OpenStudy (anonymous):

R = 9/2